### v1

@@ 3.3 Variance penalty for MMD loss
Equation Eq. 24 shows that the ensemble transport filter with a variance penalty MMD loss has a similar form compared to the EnKF with a linear assumption and linear kernel. Denote
$\tilde{\boldsymbol{X}}_{k}=\hat{\boldsymbol{X}}_{k+1}-\mathbb{E}[\hat{\boldsymbol{X}}_{k+1}],\qquad\tilde{\boldsymbol{Y}}_{k}=\mathcal{H}(\hat{\boldsymbol{X}}_{k+1})-\mathbb{E}\left[\mathcal{H}(\hat{\boldsymbol{X}}_{k+1})\right].$
@@
(28) $\mathbf{K}={\mathbb{E}\left[\tilde{\boldsymbol{X}}_{k}\tilde{\boldsymbol{Y}}_{k}^{\top}\right]}\left\{\mathbb{E}\left[\tilde{\boldsymbol{Y}}_{k}\tilde{\boldsymbol{Y}}_{k}^{\top}\right]+\mathbb{E}[\boldsymbol{\varepsilon}_{k+1}\boldsymbol{\varepsilon}_{k+1}^{\top}]\right\}^{-1}.$
The main difference between EnTranF and EnKF is that the means of $\hat{\boldsymbol{X}}_{k+1}$ and $\mathcal{H}(\hat{\boldsymbol{X}}_{k+1})$ are approximated by $\mathbb{E}[{\boldsymbol{X}}_{k+1}]$ and $\boldsymbol{y}_{k}$, respectively, in EnTranF. This drives the prior particles towards the posterior mean as dictated by transport map Eq. 23.
@@
When the RKHS kernel takes $k(\boldsymbol{x},\boldsymbol{x}^{\prime})=\boldsymbol{x}^{\top}\boldsymbol{x}^{\prime}+1$ and take regularization parameter to be $\lambda=0$, we follow the proof of Theorem 3.6 and get
(29) $\mathbf{T}=\mathbb{E}[\bar{\boldsymbol{X}}_{k}]\mathbb{E}\left[\bar{\boldsymbol{Y}}_{k}\right]^{\top}\left\{\mathbb{E}\left[\bar{\boldsymbol{Y}}_{k}\right]\mathbb{E}\left[\bar{\boldsymbol{Y}}_{k}\right]^{\top}\right\}^{-1}.$
Formally, Eq. 24 provides a complete slope in the statistical least squares sense, while Eq. 29 gives the geometric slope from the prior mean to the posterior mean. However, when the number of particles is small, Eq. 29 may be numerically unstable.
@@ 4 Numerical Results
In this section, we will use a few numerical examples to illustrate the advantage of the EnTranF over ensemble Kalman filtering in the data assimilation. A static inverse problem is implemented to show the capability of EnTranF in approximating posterior distributions in Section 4.1. In Section 4.2, we discuss the performance of EnTranF in nonlinear system by tracking Double-Well system. For strongly nonlinear cases, we consider the filtering problem of the Lorenz63 system with partial observations in Section 4.3. Finally, the performance of EnTranF for tracking high-dimensional systems is demonstrated in Section 4.4.
@@
$\text{RMSE}_{k}:=\|\bar{\boldsymbol{x}}_{k}-\boldsymbol{x}_{k}^{*}\|_{2}/\sqrt{n}$
at any assimilation step $k$. Here $\bar{\boldsymbol{x}}_{k}\in\mathbb{R}^{n}$ is the (ensemble) mean state generated by the data assimilation method, and $\boldsymbol{x}_{k}^{*}\in\mathbb{R}^{n}$ is the reference state. Considering all assimilation windows throughout the filtering process, we compute the time-averaged RMSE. Meanwhile, to eliminate the influence of randomness, we define the final performance metric RMSE as its expectation, i.e.,
(30) $\text{RMSE}=\mathbb{E}\left[\frac{1}{T}\sum\limits_{k=1}^{T}\|\bar{\boldsymbol{x}}_{k}-\boldsymbol{x}_{k}^{*}\|_{2}/\sqrt{n}\right].$
@@
(31) $\text{ENS}=\mathbb{E}\left[\frac{1}{T}\sum\limits_{k=1}^{T}\left[\mathop{\text{tr}}(\mathbf{C}_{k}^{a})/n\right]^{1/2}\right].$
@@
(32) $\text{CP}=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{1}{T}\sum_{k=1}^{T}\mathbb{I}_{\left|\bar{\boldsymbol{x}}_{k,i}-\boldsymbol{x}^{*}_{k,i}\right|/\sqrt{\mathbf{C}_{k,(i,i)}^{a}}\leq t_{\alpha}}\right)\right],$
where $\mathbb{I}$ is indicator function and $t_{\alpha}$ is $\alpha$ quantiles of standard normal distribution. We computed the expectation in Eq. 30-Eq. 32 using 20 repeated tests.
@@ 4.1 Static inverse problem
Start with a one-dimensional example, and suppose $X$ is a random variable with Gaussian prior and Gaussian observation noise. The nonlinear observation operator is considered as $\mathcal{H}(x)=2x^{3}+x$. The task is to estimate the Bayesian posterior of the $X$ given values of observations $y^{o}$. In EnTranF and EnKF, particles are generated to approximate the posterior distribution. Then, a kernel density estimate method is used to evaluate the density function of ensembles. For reference, based on Bayes’ rule
@@
Figure 3: Posterior approximation of EnTranF with Linear transport map (Panel (a)-(c)) and Nonlinear transport map (Panel (d)-(f)) when taking different bandwidths of MMD loss. The example considered the influence of bandwidth at different scales ($10^{-3}$, $10^{0}$ and $10^{3}$) on EnTranF.
We select Gaussian kernel for MMD loss with three different bandwidth values $10^{-3}$, $10^{0}$, and $10^{3}$. Linear and Nonlinear transport maps are used to approximate posterior in EnTranF, and the results are shown in Fig. 3. Larger bandwidth $\text{bw}=10^{3}$ leads to more accurate estimates of the mean compared to EnKF, but it fails to capture the morphology of the posterior distribution. When a smaller bandwidth ($\text{bw}=10^{0}$ and $\text{bw}=10^{-3}$) is given, estimation of the posterior distribution is more concentrated in the high-probability region of the true posterior distribution. EnTranF becomes more sensitive to the local details of the posterior distribution when a smaller bandwidth is chosen.
The nonlinear observation operator leads to the non-Gaussian of the posterior distribution. Both EnKF and EnTranF with linear transport maps fail to approximate the posterior distribution. As shown in Fig. 3, EnTranF achieves a significant improvement in approximating posterior distribution with a nonlinear transport map. This could be caused by the substantially different sensitivity of observations to the model state for each ensemble particle in the nonlinear assimilation problem. A linear form of transport map may not well approximate a non-Gaussian posterior.
At the same time, we will examine the performance of EnTranF, which optimizes the MMD with a variance penalty in approximating the posterior distribution. The penalted EnTranF is compared to the case without a penalty term on a two-dimensional state space $\boldsymbol{x}=[\boldsymbol{x}_{1},\boldsymbol{x}_{2}]$. We still assume Gaussian prior and the nonlinear observation function $\mathcal{H}(\boldsymbol{x})=\boldsymbol{x}_{1}^{3}+\boldsymbol{x}_{2}$.
@@
Figure 4: 2-d Histogram of ensemble particles generated by EnKF, EnTranF and EnTranFp. Red ”x” represents the MAP estimation of the corresponding method.
From Fig. 4, EnTranF and EnTranFp achieve accurate estimate of the posterior distribution, while EnKF has a significant bias. The particles of EnTranFp are more concentrated than EnTranF, and the variance penalty effectively improves the approximation of the posterior distribution in high-probability regions. Table 1 lists the posterior mean and maximum a posteriori estimation(MAP) for different filtering methods. EnTranF performs the best on the posterior mean. EnTranFp performs slightly worse than EnTranF on the posterior mean, but they give a better MAP.
Table 1: Estimation posterior mean and maximum a posteriori of the 2d static inverse problem
Reference EnKF EnTranF EnTranFp
Mean $x_{0}$ 0.2434 0.2911 0.2504 0.2921
$x_{1}$ 0.5816 0.4478 0.5551 0.5565
MAP $x_{0}$ 0.5930 0.7940 0.3719 0.4925
$x_{1}$ 0.5729 0.5729 0.6533 0.5528
@@ 4.2 Double-Well potential
(33) $\mathrm{d}x$ $=(x-x^{3})\mathrm{d}t+\gamma{\mathrm{d}W_{t}},$
(34) $z$ $=0.1x^{2}+\sin(x)+\epsilon,$
where $W_{t}$ is a standard brownian motion and $\epsilon$ is Gaussian noise with $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We take $\gamma=0.8$ and $\sigma=0.5$. In this example, the fourth-order Runge-Kutta method is used to simulate the ODE system Eq. 33 with a constant discrete step size $\Delta t=0.01$. Results of EnTranF with linear and nonlinear transport map are demonstrated.
@@
Fig. 5 compares the average RMSE between EnTranF and EnKF with different observation intervals and ensemble sizes.
@@
Figure 5: Average RMSE of Double-Well system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
The left panel of Fig. 5 indicates that EnTranF exhibits significant improvement on average RMSE compared to EnKF at larger observation intervals $\Delta_{ob}>0.2$ while performing similarly at smaller observation intervals. The nonlinearity of the filtering problem increases as the observation interval increases, leading to a larger difference in RMSE between EnKF and EnTranF. EnTranF(N-G) demonstrates the best performance. However, compared to EnTranF(N-L) and EnTranF(L-G), the difference on average RMSE is small. The right panel of Fig. 5 shows that EnTranF achieves a smaller average RMSE compared to EnKF as the increase of ensemble size. Nonlinear transport maps exhibit better performance on average RMSE than linear transport maps. However, the difference is insignificant when taking the Gaussian kernel function. The EnTranF based on Gaussian kernel function keeps sufficient posterior information in filtering.
@@
Since the DoubleWell potential represents a low-dimensional system, employing a particle filter with 10,000 particles as a benchmark posterior allows us to illustrate the approximate performance of EnTranF on the true posterior filtering distribution. To this end, we assess the tracking performance of the EnTranF using metrics identical to those employed to evaluate the estimation of the true state. Results are shown in Fig. 6.
@@
Figure 6: Take posterior mean of PF with 10,000 particles as reference, Average RMSE of Double-Well system for ensemble size $N=400$ as a function of observation interval (left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (rightl).
@@
When measuring the performance of EnTranF in tracking the posterior mean, it exhibits a significant improvement on average RMSE compared to EnKF. From the left panel of Fig. 6, it can be observed that EnTranF shows an average improvement of $38.16\%$ compared to EnKF across various observation intervals. At the same time, as the observation interval increases, the rate of increase in RMSE for EnTranF is lower than that for EnKF. This implies that EnTranF has better robustness as the nonlinearity increases. EnTranF(N-L) exhibits poor performance when the observation interval is 0.1, which may stem from inadequate optimization of the nonlinear transport map and getting trapped in local minima. As shown in the right panel of Fig. 6, EnTranF gradually approaches reference solution as the number of particles increases. However, the average RMSE of EnKF do not decrease when the ensemble size becomes larger than $200$.
@@
Figure 7: Left Panel: Spread of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size. Right panel: Average coverage probability of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size
@@
While EnTranF demonstrates excellent performance in state estimation, its performance on ensemble spread is poor, leading to inadequate particle dispersion in ensemble filtering methods (see Fig. 7). EnTranF also suffers from numerical overflow when the number of particles is low. After incorporating a variance penalty term, EnTranF can effectively address these issues. Fig. 8 shows the RMSE and Spread of EnTranFp for different observation intervals and ensemble sizes.
@@
Figure 8: EnTranF with variance penalty: Average RMSE of Double-Well system for ensemble size $N=400$ as a function of observation interval (Panel (a)), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Panel (b)). Spread of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Panel (c)). Average coverage probability of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Panel (d))
@@
From Fig. 5 and Fig. 8 (a)-(b), the average RMSE of EnTranFp remains consistent across different observation intervals and ensemble sizes. The EnTranFp maintains the accurancy of EnTranF in state estimation. EnTranFp(N-L), EnTranFp(L-G) and EnTranFp(N-G) show a small improvement compare to EnTranF in average RMSE. In panel (c) and (d) of Fig. 8, EnTranFp with linear transport map exhibits a similar ensemble spread to EnKF. But it has a significantly higher probability of covering the true state. EnTranFp with a nonlinear transport map reduces the Spread while simultaneously increasing the probability of covering the true state. Compare with the results of EnTranF in Fig. 7, the EnTranFp enhances the robustness of the EnTranF and leads to a better approximation of the posterior.

### v2

@@ 3.2.3 Connection with Ensemble Kalman Filter
Equation (3.24) shows that the ensemble transport filter with a variance penalty MMD loss has a similar form compared to the EnKF with a linear assumption and linear kernel. Denote
$\tilde{\bm{X}}_{k}=\hat{\bm{X}}_{k+1}-\mathbb{E}[\hat{\bm{X}}_{k+1}],\qquad\tilde{\bm{Y}}_{k}=\mathbf{H}(\hat{\bm{X}}_{k+1})-\mathbb{E}\left[\mathbf{H}(\hat{\bm{X}}_{k+1})\right].$
@@
$\mathbf{K}={\mathbb{E}\left[\tilde{\bm{X}}_{k}\tilde{\bm{Y}}_{k}^{\top}\right]}\left\{\mathbb{E}\left[\tilde{\bm{Y}}_{k}\tilde{\bm{Y}}_{k}^{\top}\right]+\mathbb{E}[\bm{\varepsilon}_{k+1}\bm{\varepsilon}_{k+1}^{\top}]\right\}^{-1}.$ (3.28)
The main difference between EnTranF and EnKF is that the means of $\hat{\bm{X}}_{k+1}$ and $\mathbf{H}(\hat{\bm{X}}_{k+1})$ are approximated by $\mathbb{E}[{\bm{X}}_{k+1}]$ and $\bm{y}_{k}$, respectively, in EnTranF. This drives the prior particles towards the posterior mean as dictated by transport map (3.23).
@@
When the RKHS kernel takes $k(\bm{x},\bm{x}^{\prime})=\bm{x}^{\top}\bm{x}^{\prime}+1$ and take regularization parameter to be $\lambda=0$, we follow the proof of 3.4 and get
$\mathbf{T}=\mathbb{E}[\bar{\bm{X}}_{k}]\mathbb{E}\left[\bar{\bm{Y}}_{k}\right]^{\top}\left\{\mathbb{E}\left[\bar{\bm{Y}}_{k}\right]\mathbb{E}\left[\bar{\bm{Y}}_{k}\right]^{\top}\right\}^{-1}.$ (3.29)
Formally, (3.24) provides a complete slope in the statistical least squares sense, while (3.29) gives the geometric slope from the prior mean to the posterior mean. However, when the number of particles is small, (3.29) may be numerically unstable in computing the matrix inversion.
@@ 4 Numerical Results
In this section, we will use a few numerical examples to illustrate the advantage of the EnTranF over ensemble Kalman filtering in the data assimilation. A static inverse problem is implemented to show the capability of EnTranF in approximating posterior distributions in 4.1. In 4.2, we discuss the performance of EnTranF in nonlinear system by tracking Double-Well system. For strongly nonlinear cases, we consider the filtering problem of the Lorenz63 system with partial observations in 4.3. Finally, the performance of EnTranF for tracking high-dimensional systems is demonstrated in 4.4.
@@
$\text{RMSE}_{k}:=\|\bar{\bm{x}}_{k}-\bm{x}_{k}^{*}\|_{2}/\sqrt{n}$
at any assimilation step $k$. Here $\bar{\bm{x}}_{k}\in\mathbb{R}^{n}$ is the (ensemble) mean state generated by the data assimilation method, and $\bm{x}_{k}^{*}\in\mathbb{R}^{n}$ is the reference state. Considering all assimilation windows throughout the filtering process, we compute the time-averaged RMSE. Meanwhile, to eliminate the influence of randomness, we define the final performance metric RMSE as its expectation, i.e.,
$\text{RMSE}=\mathbb{E}\left[\frac{1}{T}\sum\limits_{k=1}^{T}\|\bar{\bm{x}}_{k}-\bm{x}_{k}^{*}\|_{2}/\sqrt{n}\right].$ (4.30)
@@
$\text{ENS}=\mathbb{E}\left[\frac{1}{T}\sum\limits_{k=1}^{T}\left[\mathop{\text{tr}}(\mathbf{C}_{k}^{a})/n\right]^{1/2}\right].$ (4.31)
@@
$\text{CP}=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{1}{T}\sum_{k=1}^{T}\mathbb{I}_{\left|\bar{\bm{x}}_{k,i}-\bm{x}^{*}_{k,i}\right|/\sqrt{\mathbf{C}_{k,(i,i)}^{a}}\leq t_{\alpha}}\right)\right],$ (4.32)
where $\mathbb{I}$ is indicator function and $t_{\alpha}$ is $\alpha$ quantiles of standard normal distribution. We take $\alpha=0.025$ and compute the expectation in (4.30)-(4.32) using 20 repeated tests.
@@ 4.1 Static inverse problem
Start with a one-dimensional example, and suppose $X$ is a random variable with Gaussian prior $\mathcal{N}(0.5,1)$ and Gaussian observation noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$. The nonlinear observation operator is considered as ${H}(x)=2x^{3}+x$. The task is to estimate the Bayesian posterior of the $X$ given values of observations $y^{o}=1.2$. In EnTranF and EnKF, particles are generated to approximate the posterior distribution. Then, a kernel density estimate method is used to evaluate the density function of ensembles. For reference, based on Bayes’ rule
@@
Figure 4.4: Posterior approximation of EnTranF with Linear transport map (Panel (a)-(c)) and Nonlinear transport map (Panel (d)-(f)) when taking different bandwidths ($^{-3}, 1001$$^{0}$ and $10^{3}$).
We select Gaussian kernel for MMD loss with three different bandwidth values $^{-3}, 1001$$^{0}$, and $10^{3}$. Linear and Nonlinear transport maps are used to approximate posterior in EnTranF, and the results are shown in Figure 4.4. Larger bandwidth $\text{bw}=10^{3}$ leads to more accurate estimates of the mean compared to EnKF, but it fails to capture the morphology of the posterior distribution. When a smaller bandwidth ($\text{bw}=10^{0}$ and $\text{bw}=10^{-3}$) is given, estimation of the posterior distribution is more concentrated in the high-probability region of the true posterior distribution. EnTranF becomes more sensitive to the local details of the posterior distribution when a smaller bandwidth is chosen.
The nonlinear observation operator leads to the non-Gaussian of the posterior distribution. Both EnKF and EnTranF with linear transport maps fail to approximate the posterior distribution. As shown in Figure 4.4, EnTranF achieves a significant improvement in approximating posterior distribution with a nonlinear transport map. This could be caused by the substantially different sensitivity of observations to the model state for each ensemble particle in the nonlinear assimilation problem. A linear form of transport map may not well approximate a non-Gaussian posterior.
At the same time, we will examine the performance of EnTranF, which optimizes the MMD with a variance penalty in approximating the posterior distribution. The penalted EnTranF is compared to the case without a penalty term on a two-dimensional state space $\bm{x}=[\bm{x}_{1},\bm{x}_{2}]$. We assume Gaussian prior $\mathcal{N}\left(\left[\begin{matrix}0.5\\ 0.5\end{matrix}\right],\left[\begin{matrix}1&0\\ 0&1\end{matrix}\right]\right)$ and the nonlinear observation $y^{o}=0.8$ with operator $\mathbf{H}(\bm{x})=\bm{x}_{1}^{3}+\bm{x}_{2}$ and additive Gaussian noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$.
@@
Figure 4.5: 2-d Histogram of ensemble particles generated by EnKF, EnTranF and EnTranFp. Red $\times$” represents the MAP estimation of the corresponding method.
From Figure 4.5, EnTranF and EnTranFp achieve accurate estimate of the posterior distribution, while EnKF has a significant bias. The particles of EnTranFp are more concentrated than EnTranF, and the variance penalty effectively improves the approximation of the posterior distribution in high-probability regions. Table 1 lists the average RMSE and ENS of different filtering methods with different ensemble sizes. The EnTranFp achieves the lowest RMSE and exhibits the highest stability across both ensemble sizes (200 and 400). EnTranF shows slightly improved performance over EnTranFp at larger ensemble sizes (800), both variants significantly outperform EnKF in terms of RMSE. As for the ENS, we use a particle filter with 10,000 particles as a reference benchmark. As shown, the ENS of EnTranF and EnTranFp are notably closer to this reference than those of the EnKF. These results highlight the effectiveness of our approach in generating more accurate posterior estimates. Furthermore, the advantage of EnTranFp shows the importance of the variance penalty in enhancing both the accuracy and stability of the transport map. Figure 4.6 shows the relationship between the distribution approximation accuracy and computational time for EnKF, EnTranF, and EnTranFp. EnKF requires less computational time but has lower distribution approximation accuracy; EnTranF, due to the need to optimize the loss function, consumes more computational time, but achieves higher distribution approximation accuracy.
Table 1: Average RMSE and ENS of EnKF, EntranF and EnTranFp for the 2d static inverse problem with different ensemble size.
Results of RMSE
EnKF EnTranF EnTranFp
200 0.1363($\pm$0.0862) 0.1377($\pm$0.0732) 0.1255($\pm$0.0592)
400 0.1543($\pm$0.0510) 0.0962($\pm$0.0554) 0.0878($\pm$0.0498)
800 0.1329($\pm$0.0381) 0.0702($\pm$0.0331) 0.0742($\pm$0.0350)
Results of ENS
Ref. EnKF EnTranF EnTranFp
200 0.6076 0.8340($\pm$0.0387) 0.8197($\pm$0.0728) 0.7616($\pm$0.0441)
400 0.6076 0.8216($\pm$0.0225) 0.8059($\pm$0.0664) 0.7466($\pm$0.0455)
800 0.6076 0.8332($\pm$0.0252) 0.7636($\pm$0.0286) 0.7306($\pm$0.0221)
Figure 4.6: log-log plot of computation time and MMD of EnKF, EnTranF and EnTranFp.
@@ 4.2 Double-Well potential
$\mathrm{d}x$ $=(x-x^{3})\mathrm{d}t+\gamma{\mathrm{d}W_{t}},$ (4.33)
$z$ $=0.1x^{2}+\sin(x)+\epsilon,$ (4.34)
where $W_{t}$ is a standard brownian motion and $\epsilon$ is Gaussian noise with $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We take $\gamma=0.8$ and $\sigma=0.5$. In this example, the fourth-order Runge-Kutta method is used to simulate the ODE system (4.33) with a constant discrete step size $\Delta t=0.01$. Results of EnTranF with linear and nonlinear transport map are demonstrated. For nonlinear transport map, a shallow DNN comprising $L=3$ layers with width $\{1,10,1\}$ and tanh activation functions is used to define the nudging term of the nonlinear transport map. The bandwidth of MMD was set to be the median of all $L_{2}$ distance between each pair of particles.
@@
Figure 4.7 compares the average RMSE between EnTranF and EnKF with different observation intervals and ensemble sizes.
@@
Figure 4.7: Average RMSE of Double-Well system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
The left panel of Figure 4.7 indicates that EnTranF exhibits significant improvement on average RMSE compared to EnKF at larger observation intervals $\Delta_{ob}>0.2$ while performing similarly at smaller observation intervals. The nonlinearity of the filtering problem increases as the observation interval increases, leading to a larger difference in RMSE between EnKF and EnTranF. EnTranF(N-G) demonstrates the best performance. However, compared to EnTranF(N-L) and EnTranF(L-G), the difference on average RMSE is small. The right panel of Figure 4.7 shows that EnTranF achieves a smaller average RMSE compared to EnKF as the increase of ensemble size. Nonlinear transport maps exhibit better performance on average RMSE than linear transport maps. However, the difference is insignificant when taking the Gaussian kernel function. The EnTranF based on Gaussian kernel function keeps sufficient posterior information in filtering.
@@
Since the DoubleWell potential represents a low-dimensional system, employing a particle filter with 10,000 particles as a benchmark posterior allows us to illustrate the approximate performance of EnTranF on the true posterior filtering distribution. To this end, we assess the tracking performance of the EnTranF using metrics identical to those employed to evaluate the estimation of the true state. Results are shown in Figure 4.8.
@@
Figure 4.8: Take posterior mean of PF with 10,000 particles as reference, Average RMSE of Double-Well system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
When measuring the performance of EnTranF in tracking the posterior mean, it exhibits a significant improvement on average RMSE compared to EnKF. From the left panel of Figure 4.8, it can be observed that EnTranF shows an average improvement of $38.16\%$ compared to EnKF across various observation intervals. At the same time, as the observation interval increases, the rate of increase in RMSE for EnTranF is lower than that for EnKF. This implies that EnTranF has better robustness as the nonlinearity increases. EnTranF(N-L) exhibits poor performance when the observation interval is 0.1, which may stem from inadequate optimization of the nonlinear transport map and getting trapped in local minima. As shown in the right panel of Figure 4.8, EnTranF gradually approaches reference solution as the number of particles increases. However, the average RMSE of EnKF do not decrease when the ensemble size becomes larger than $200$.
@@
Figure 4.9: Left Panel: Spread of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size. Right panel: Average coverage probability of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size
@@
While EnTranF demonstrates excellent performance in state estimation, its performance on ensemble spread is poor, leading to inadequate particle dispersion in ensemble filtering methods (see Figure 4.9). EnTranF also suffers from numerical overflow when the number of particles is low. After incorporating a variance penalty term, EnTranF can effectively address these issues. Figure 4.10 shows the RMSE and Spread of EnTranFp for different observation intervals and ensemble sizes.
@@
Figure 4.10: EnTranF with variance penalty: Average RMSE of Double-Well system for ensemble size $N=400$ as a function of observation interval (Panel (a)), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Panel (b)). Spread of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Panel (c)). Average coverage probability of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Panel (d))
@@
From Figure 4.7 and Figure 4.10 (a)-(b), the average RMSE of EnTranFp remains consistent across different observation intervals and ensemble sizes. The EnTranFp maintains the accurancy of EnTranF in state estimation. EnTranFp(N-L), EnTranFp(L-G) and EnTranFp(N-G) show a small improvement compare to EnTranF in average RMSE. In panel (c) and (d) of Figure 4.10, EnTranFp with linear transport map exhibits a similar ensemble spread to EnKF. But it has a significantly higher probability of covering the true state. EnTranFp with a nonlinear transport map reduces the Spread while simultaneously increasing the probability of covering the true state. Compare with the results of EnTranF in Figure 4.9, the EnTranFp enhances the robustness of the EnTranF and leads to a better approximation of the posterior. Finally, we directly compared the errors in the posterior variances estimated by EnTranF and EnKF against the posterior variance estimated by a particle filter with 10,000 particles used as a reference, as shown in Figure 4.11. EnTranFp (L-G) and EnTranFp (N-G) get better accurate variance estimates compared to EnKF, while EnTranFp with linear kernel get a poor estimate of variance.
Figure 4.11: EnTranF with variance penalty: Average covariance error of Double-Well system for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size.